what is the decimal answer for the fraction 7 over 20
1
Expert's answer
2013-02-04T10:34:50-0500
To obtain decimal answer we need to turn given fraction into one with denominator being power of ten. Let's do that: as we have 20 in the denominator it's easy to obtain 100 by multiplying both numerator and denominator by 57/20=(5*7)/(5*20)=35/100=0.35so the answer is 0.35